OpenStudy (anonymous):

Steve is scheduled to work 5/6 of an hour at the school fair he has already worked 2/3 of an hour how much longer does he have to work?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Wait out of these which would it be. 5/6? 1h? 2/3? 1/6?

OpenStudy (anonymous):

2/3+x=5/6 x = (5/6)(-2/3) x = (5-4)/6 x= 1/6 If I deserve medal then give me one

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

thank

OpenStudy (anonymous):

One more plz, if thats ok XD

OpenStudy (anonymous):

I try

OpenStudy (anonymous):

Thanks you are making scarves for presents each scarf needs 3/4yd of fabric how many yards of fabric do you need for 7 scarves?

OpenStudy (anonymous):

if you have another post you can come back I'm slow

OpenStudy (anonymous):

9 and 1/3 5 and 1/4 10 and 2/3 6yd?

OpenStudy (anonymous):

ok! thanks

OpenStudy (anonymous):

(3/4)*7=21/4 5and 1/4

OpenStudy (anonymous):

Thank you so much! I think im going to get a A on my test cuz of you XD

OpenStudy (anonymous):

you welcome, come back if you need help

OpenStudy (anonymous):

Thanks